Ratios are mostly calculated using data drawn from the financial statements of a firm. However, another group of ratios, called market-based ratios, relate to a firm’s observable market value, stock prices, and book values, integrating information from both the market and the firm’s financial statements.
Consider the case of Green Caterpillar Garden Supplies Inc.:
Green Caterpillar Garden Supplies Inc. just reported earnings after tax (also called net income) of $95,000,000, and a current stock price of $28.50 per share. The company is forecasting an increase of 25% for its after-tax income next year, but it also expects it will have to issue 2,800,000 new shares of stock (raising its shares outstanding from 5,500,000 to 8,300,000).
If Green Caterpillar’s forecast turns out to be correct and its price-to-earnings (P/E) ratio does not change, what does the company’s management expect its stock price to be one year from now? (Note: Round intermediate calculations to four decimal places. Round the expected stock price to two decimal places.)
$23.61 per share
$28 per share
$17.71 per share
$29.51 per share

One year later, Green Caterpillar’s shares are trading at $54.56 per share, and the company reports the value of its total common equity as $39,192,600. Given this information, Green Caterpillar’s market-to-book (M/B) ratio is (Note: Do not round intermediate calculations.)
Is it possible for a company to exhibit a negative EPS and thus a negative P/E ratio?
No
Yes
Which of the following statements is true about market value ratios?
Companies with high research and development (R&D) expenses tend to have low P/E ratios.
Companies with high research and development (R&D) expenses tend to have high P/E ratios.
Companies have the opportunity to use varying amounts of different sources of financing to acquire their assets, including internal and external sources, and debt (borrowed) and equity funds.

Company A uses long-term debt to finance its assets, and company B uses capital generated from shareholders to finance its assets. Which company would be considered a financially leveraged firm?
Company B
Company A

Which of the following is true about the leveraging effect?
Interest on debt can be deducted from pre-tax income, resulting in a greater taxable income and a smaller available operating income.
Interest on debt is a tax deductible expense, which means that it can reduce a firm’s taxable income and tax obligation.
Blue Sky Drone Company has a total asset turnover ratio of 6.00x, net annual sales of $25 million, and operating expenses of $11.25 million (including depreciation and amortization). On its current balance sheet and income statement, respectively, it reported total debt of $2.5 million, on which it pays 11% interest on its outstanding debt.
To analyze a company’s financial leverage situation, you need to measure the firm’s debt management ratios. Based on the preceding information, what are the values for Blue Sky Drone’s debt management ratios? (Note: Do not round intermediate calculations.)
Ratio
Value
Debt ratio
Options 60.00 / 48.00 / 138.00 / 78.00
Times-interest-earned ratio
Options 90.00 / 37.50 / 50.00 / 25.00

Blue Sky Drone Company raises around ____ from creditors for each dollar of equity. Options 60.00 / 1.80 / 1.50 / 2.10

Influenced by a firm’s ability to make interest payments and pay back its debt, if all else is equal, creditors would prefer to give loans to companies with ___ times-interest-earned ratios (TIE). Options low / high
Answer #1

a). Net income = 95,000,000; shares O/S = 5,500,000

EPS = 95,000,000/5,500,000 = 17.27

P/E ratio = share price/EPS = 28.50/1.17.27 = 1.65

Net income next year = 95,000,000*(1+25%) = 118,750,000

Shares O/S next year = 8,300,000

EPS next year = 118,750,000/8,300,000 = 14.31

If P/E ratio remains at 1.65 then price next year = 1.65*14.31 = $23.61 per share

b). Share price = 54.56; shares O/S = 8,300,000

Market value of equity = 54.56*8,300,000 = 452,848,000

Book value of equity = 39,192,600

Market to book ratio = 452,848,000/39,192,600 = 11.55

c). Yes, it is possible for a company to have negative EPS when a company is showing net loss, instead of profit. In that case, P/E ratio will be negative but in such cases, P/E is not reported as negative P/E does not give any information about the company. It is usually shown as "not applicable".

d). Companies with high research and development (R&D) expenses tend to have high P/E ratios, as earnings will be low due to the high R&D expenses, leading to a high P/E ratio.

e). Company A would be considered a financially leveraged firm as it is using debt as a source of capital.

f). Interest on debt is a tax deductible expense, which means that it can reduce a firm’s taxable income and tax obligation.

g). Total asset turnover = Net sales/Total assets = 6

Total assets = 25/6 = 4.167 million

Debt = 2.5 million

Debt ratio = debt/total assets = 2.5/4.167 = 60%

Times interest earned = EBIT/interest expense

EBIT = Net sales-operating expenses = 25-11.25 = 13.75 million

Interest expense = 11%*2.5 = 0.275 million

Times interest earned = 13.75/0.275 = 50.00

Blue Sky Drone Company is raising around $1.50 from creditors for each dollar of equity. (Debt to equity = 60/40 = 1.5)

Creditors would prefer to give loans to companies with high TIE as it indicates that the company can service the debt comfortably.